OpenStudy (anonymous):

There are 6 red apples, 4 yellow apples, and 2 green apples in a bucket. Maria will chose two apples at random without replacement. What is the probability that Maria will choose a red apple and a green apple?

OpenStudy (anonymous):

thank you but unfourtanettly that is not one of the answer choices

OpenStudy (anonymous):

1/11 ?

OpenStudy (anonymous):

yea thats one

OpenStudy (anonymous):

\[ \huge \frac {\binom 6 1 \times \binom 2 1}{12\times 11} = \frac 1 {11} \] I should have read the question carefully.

OpenStudy (anonymous):

thank you so very much